OpenStudy (anonymous):

in the system -12c+3d=6and-3c+6d=-9 ?

OpenStudy (anonymous):

multiply .. 2 with first eq ..... d term will get cancel u will get the c value substitute the c value in anyone eq to get d ..

OpenStudy (anonymous):

hope u get that ... @freeman1

OpenStudy (anonymous):

2*(-12c+3d=6) -24c+6d=12 (-)-3c+6d=-9 ------------- -21c =21 c=-1 -12c+3d=6 -12(-1)+3d=6 12+3d=6 3d=-6 d=-2 done >>